OpenStudy (anonymous):

Solve the equation 5/13t=-9 ( 5/13 is a fraction)

OpenStudy (anonymous):

\[\frac{ 5t }{ 13 }=-9\]

OpenStudy (anonymous):

Cross Multiply

OpenStudy (anonymous):

we need to get rid of the fraction so that x is by itself

OpenStudy (anonymous):

if we multiply each side by the reciprocal the fraction of the left will disappear because it will be 1

OpenStudy (anonymous):

5t = -9*13 5t=-117 t=-117/5

OpenStudy (anonymous):

\[\frac{ 13 }{ 5 }\times \frac{ 5t }{ 13 }=-9\times \frac{ 13 }{ 5 }\]